I need help solving the following question with a calculator and how to type it in
The heights of adult men in America are normally distributed, with a mean of 69.2 inches and a standard deviation of 2.68 inches. The heights of adult women in America are also normally distributed, but with a mean of 64.4 inches and a standard deviation of 2.59 inches.
a) If a man is 6 feet 3 inches tall, what is his z-score (to two decimal places)?
z =
b) If a woman is 5 feet 11 inches tall, what is her z-score (to two decimal places)?
z =
c) Who is relatively taller?

Answer by Theo(13342)

the mean for the men is 69.2 and the standard deviation is 2.68.

the mean for the women is 64.4 and the standard deviation is 2.59.

a man is 6 foot 3 = 75 inches.

the z-score for the man is z = (75 - 69.2) / 2.68 = 2.164179104.

a woman is 5 foot 11 = 71 inches.

the z-score for the woman is z = (71 - 64.4) / 2.59 = 2.548262548.

relative to their respective populations, the woman is taller than the man.

if you look up their z-scores to find the area to the left of each, the man's z-score is greater than 98.47747514% of the male population and the woman's z-score is greater than 99.45869278% of the female population.

i used the TI-84 Plus to get these figures.

an online calculator that is relatively easy to use can be found at http://davidmlane.com/hyperstat/z_table.html

this calculator can work off the z-score or off the raw score.

when working off the z-score, the mean is 0 and the standard deviation is 1.

when working off the z-score,the mean is the population mean and the standard deviation is the population standard deviation.

using this calculator, i get the following results.

for the men.

$$$

$$$

for the women.

$$$

$$$